# EducationQ: Evaluating LLMs’ Teaching Capabilities Through Multi-Agent Dialogue Framework

### v2

@@ 5 Evaluation Metrics
    $PNIR=\frac{N_{neg}}{N_{pos}}\vspace{-0.2cm}$ (2)
@@
    $CSS=\sigma(SLGPD)\vspace{-0.3cm}$ (3)
@@ 1 MMLU-Pro
Business $100$ $6.68$
2 Law $100$ $6.68$
3 Psychology $100$ $6.68$
4 Biology $100$ $6.68$
5 Chemistry $100$ $6.68$
6 History $100$ $6.68$
7 Health $100$ $6.68$ 86.78
8 Economics $100$ $6.68$
9 Math $100$ $6.68$
10 Physics $100$ $6.68$
11 Engineering $100$ $6.68$
12 Philosophy $100$ $6.68$
13 Computer Science $100$ $6.68$
@@ 14 GPQA
Physics $86$ $5.74$
15 Chemistry $93$ $6.21$ 13.22
16 Biology $19$ $1.27$
@@ Appendix E Teaching Performance Ranking by Subject
1 Gemini 1.5 Pro 002 $8.00$
2 OpenAI o1-mini $5.00$
2 Claude 3.5 Sonnet $5.00$
2 Llama 3.1 70B Instruct $5.00$
5 WizardLM-2 8x22B $4.00$
6 Llama 3.1 405B Instruct $3.00$
6 GPT-4o-mini $3.00$
8 Llama 3.1 8B Instruct $2.00$
8 Mistral Nemo $2.00$
8 DeepSeek V2.5 $2.00$
8 Phi-3.5-mini Instruct $2.00$
12 Qwen 2.5 72B Instruct $0.00$
12 Command R 08-2024 $0.00$
14 Hermes 3 Llama 3.1 70B $-1.00$
@@
1 Llama 3.1 70B Instruct $11.00$
2 Qwen 2.5 72B Instruct $7.00$
3 Gemini 1.5 Pro 002 $6.00$
4 Claude 3.5 Sonnet $2.00$
4 Hermes 3 Llama 3.1 70B $2.00$
4 WizardLM-2 8x22B $2.00$
7 Llama 3.1 8B Instruct $1.00$
8 OpenAI o1-mini $0.00$
8 GPT-4o-mini $0.00$
8 DeepSeek V2.5 $0.00$
8 Command R 08-2024 $0.00$
12 Mistral Nemo $-1.00$
13 Llama 3.1 405B Instruct $-4.00$
13 Phi-3.5-mini Instruct $-4.00$
@@
1 Llama 3.1 70B Instruct $18.00$
2 Gemini 1.5 Pro 002 $12.00$
3 Llama 3.1 405B Instruct $10.00$
4 Llama 3.1 8B Instruct $9.00$
5 Command R 08-2024 $7.00$
6 OpenAI o1-mini $6.00$
6 Hermes 3 Llama 3.1 70B $6.00$
6 Mistral Nemo $6.00$
9 GPT-4o-mini $4.00$
9 DeepSeek V2.5 $4.00$
11 Claude 3.5 Sonnet $3.00$
11 Qwen 2.5 72B Instruct $3.00$
11 WizardLM-2 8x22B $3.00$
14 Phi-3.5-mini Instruct $1.00$
@@
1 Llama 3.1 70B Instruct $10.08$
2 Gemini 1.5 Pro 002 $9.24$
3 Qwen 2.5 72B Instruct $5.88$
4 Llama 3.1 405B Instruct $3.36$
4 Llama 3.1 8B Instruct $3.36$
6 Hermes 3 Llama 3.1 70B $2.52$
7 OpenAI o1-mini $1.68$
7 DeepSeek V2.5 $1.68$
7 Mistral Nemo $1.68$
10 Command R 08-2024 $0.84$
10 Phi-3.5-mini Instruct $0.84$
10 WizardLM-2 8x22B $0.84$
13 Claude 3.5 Sonnet $0.00$
13 GPT-4o-mini $0.00$
@@
1 Llama 3.1 70B Instruct $12.44$
2 Qwen 2.5 72B Instruct $8.81$
3 OpenAI o1-mini $8.29$
4 Gemini 1.5 Pro 002 $7.77$
5 Claude 3.5 Sonnet $6.74$
5 Mistral Nemo $6.74$
7 Llama 3.1 405B Instruct $6.22$
7 GPT-4o-mini $6.22$
9 DeepSeek V2.5 $4.15$
9 Llama 3.1 8B Instruct $4.15$
9 Command R 08-2024 $4.15$
9 Phi-3.5-mini Instruct $4.15$
9 WizardLM-2 8x22B $4.15$
14 Hermes 3 Llama 3.1 70B $3.63$
@@
1 Llama 3.1 70B Instruct $14.00$
2 Llama 3.1 8B Instruct $10.00$
3 Llama 3.1 405B Instruct $8.00$
4 Hermes 3 Llama 3.1 70B $3.00$
5 OpenAI o1-mini $2.02$
6 Gemini 1.5 Pro 002 $2.00$
6 Mistral Nemo $2.00$
8 Qwen 2.5 72B Instruct $1.00$
8 Command R 08-2024 $1.00$
8 WizardLM-2 8x22B $1.00$
11 Claude 3.5 Sonnet $0.00$
11 DeepSeek V2.5 $0.00$
13 Phi-3.5-mini Instruct $-1.00$
14 GPT-4o-mini $-1.01$
@@
1 Llama 3.1 70B Instruct $24.00$
2 Gemini 1.5 Pro 002 $13.00$
3 Llama 3.1 405B Instruct $11.00$
4 Qwen 2.5 72B Instruct $6.00$
4 Mistral Nemo $6.00$
6 OpenAI o1-mini $5.00$
7 Llama 3.1 8B Instruct $4.00$
7 Hermes 3 Llama 3.1 70B $4.00$
7 WizardLM-2 8x22B $4.00$
10 DeepSeek V2.5 $3.00$
10 Command R 08-2024 $3.00$
12 Claude 3.5 Sonnet $2.00$
12 Phi-3.5-mini Instruct $2.00$
14 GPT-4o-mini $1.00$
@@
1 Gemini 1.5 Pro 002 $9.00$
1 Llama 3.1 70B Instruct $9.00$
3 Llama 3.1 8B Instruct $8.00$
3 Mistral Nemo $8.00$
5 Command R 08-2024 $6.00$
5 WizardLM-2 8x22B $6.00$
7 OpenAI o1-mini $5.00$
7 Claude 3.5 Sonnet $5.00$
7 Hermes 3 Llama 3.1 70B $5.00$
7 DeepSeek V2.5 $5.00$
11 Llama 3.1 405B Instruct $4.00$
11 Qwen 2.5 72B Instruct $4.00$
11 GPT-4o-mini $4.00$
14 Phi-3.5-mini Instruct $1.00$
@@
1 OpenAI o1-mini $9.00$
2 Llama 3.1 405B Instruct $8.00$
3 Claude 3.5 Sonnet $7.00$
3 Gemini 1.5 Pro 002 $7.00$
3 WizardLM-2 8x22B $7.00$
6 Llama 3.1 8B Instruct $4.00$
6 Hermes 3 Llama 3.1 70B $4.00$
8 Qwen 2.5 72B Instruct $3.00$
8 Mistral Nemo $3.00$
10 GPT-4o-mini $2.00$
10 Llama 3.1 70B Instruct $2.00$
10 DeepSeek V2.5 $2.00$
10 Command R 08-2024 $2.00$
14 Phi-3.5-mini Instruct $-4.00$
@@
1 OpenAI o1-mini $8.60$
2 Qwen 2.5 72B Instruct $6.99$
3 Llama 3.1 70B Instruct $6.45$
4 Llama 3.1 405B Instruct $5.38$
5 Hermes 3 Llama 3.1 70B $4.84$
6 Gemini 1.5 Pro 002 $4.30$
6 Llama 3.1 8B Instruct $4.30$
8 Claude 3.5 Sonnet $3.76$
9 Mistral Nemo $1.61$
9 DeepSeek V2.5 $1.61$
9 Command R 08-2024 $1.61$
12 GPT-4o-mini $1.08$
13 Phi-3.5-mini Instruct $0.54$
14 WizardLM-2 8x22B $0.00$
@@
1 Llama 3.1 70B Instruct $8.00$
1 Llama 3.1 405B Instruct $8.00$
3 WizardLM-2 8x22B $7.00$
4 OpenAI o1-mini $6.00$
4 Qwen 2.5 72B Instruct $6.00$
6 Gemini 1.5 Pro 002 $4.00$
6 GPT-4o-mini $4.00$
6 DeepSeek V2.5 $4.00$
6 Phi-3.5-mini Instruct $4.00$
10 Claude 3.5 Sonnet $3.00$
11 Command R 08-2024 $2.00$
11 Mistral Nemo $2.00$
13 Hermes 3 Llama 3.1 70B $1.00$
14 Llama 3.1 8B Instruct $-1.00$
@@
1 Llama 3.1 70B Instruct $19.00$
2 Llama 3.1 8B Instruct $12.00$
3 Gemini 1.5 Pro 002 $10.00$
3 Llama 3.1 405B Instruct $10.00$
5 Hermes 3 Llama 3.1 70B $9.00$
6 OpenAI o1-mini $8.00$
7 Mistral Nemo $7.00$
8 Claude 3.5 Sonnet $6.00$
8 WizardLM-2 8x22B $6.00$
10 DeepSeek V2.5 $5.00$
11 Qwen 2.5 72B Instruct $4.00$
11 Command R 08-2024 $4.00$
13 GPT-4o-mini $2.00$
14 Phi-3.5-mini Instruct $1.00$
@@
1 Hermes 3 Llama 3.1 70B $10.00$
2 OpenAI o1-mini $9.00$
2 DeepSeek V2.5 $9.00$
4 Gemini 1.5 Pro 002 $8.00$
4 Llama 3.1 405B Instruct $8.00$
4 Qwen 2.5 72B Instruct $8.00$
7 Llama 3.1 70B Instruct $6.00$
8 WizardLM-2 8x22B $5.00$
8 Command R 08-2024 $5.00$
10 Llama 3.1 8B Instruct $4.00$
10 Mistral Nemo $4.00$
10 Claude 3.5 Sonnet $4.00$
10 Phi-3.5-mini Instruct $4.00$
14 GPT-4o-mini $3.00$

### v3

@@ 5 Evaluation Metrics
    $PNIR=\frac{N_{neg}}{N_{pos}}\vskip-5.69046pt$ (2)
@@
    $CSS=\sigma(SLGPD)\vskip-8.5359pt$ (3)
@@ 1 MMLU-Pro
Business 100100100 6.686.686.68
2 Law 100100100 6.686.686.68
3 Psychology 100100100 6.686.686.68
4 Biology 100100100 6.686.686.68
5 Chemistry 100100100 6.686.686.68
6 History 100100100 6.686.686.68
7 Health 100100100 6.686.686.68 86.78
8 Economics 100100100 6.686.686.68
9 Math 100100100 6.686.686.68
10 Physics 100100100 6.686.686.68
11 Engineering 100100100 6.686.686.68
12 Philosophy 100100100 6.686.686.68
13 Computer Science 100100100 6.686.686.68
@@ 14 GPQA
Physics 868686 5.745.745.74
15 Chemistry 939393 6.216.216.21 13.22
16 Biology 191919 1.271.271.27
@@ Appendix E Teaching Performance Ranking by Subject
1 Gemini 1.5 Pro 002 8.008.008.00
2 OpenAI o1-mini 5.005.005.00
2 Claude 3.5 Sonnet 5.005.005.00
2 Llama 3.1 70B Instruct 5.005.005.00
5 WizardLM-2 8x22B 4.004.004.00
6 Llama 3.1 405B Instruct 3.003.003.00
6 GPT-4o-mini 3.003.003.00
8 Llama 3.1 8B Instruct 2.002.002.00
8 Mistral Nemo 2.002.002.00
8 DeepSeek V2.5 2.002.002.00
8 Phi-3.5-mini Instruct 2.002.002.00
12 Qwen 2.5 72B Instruct 0.000.000.00
12 Command R 08-2024 0.000.000.00
14 Hermes 3 Llama 3.1 70B 1.00-1.00- 1.00
@@
1 Llama 3.1 70B Instruct 11.0011.0011.00
2 Qwen 2.5 72B Instruct 7.007.007.00
3 Gemini 1.5 Pro 002 6.006.006.00
4 Claude 3.5 Sonnet 2.002.002.00
4 Hermes 3 Llama 3.1 70B 2.002.002.00
4 WizardLM-2 8x22B 2.002.002.00
7 Llama 3.1 8B Instruct 1.001.001.00
8 OpenAI o1-mini 0.000.000.00
8 GPT-4o-mini 0.000.000.00
8 DeepSeek V2.5 0.000.000.00
8 Command R 08-2024 0.000.000.00
12 Mistral Nemo 1.00-1.00- 1.00
13 Llama 3.1 405B Instruct 4.00-4.00- 4.00
13 Phi-3.5-mini Instruct 4.00-4.00- 4.00
@@
1 Llama 3.1 70B Instruct 18.0018.0018.00
2 Gemini 1.5 Pro 002 12.0012.0012.00
3 Llama 3.1 405B Instruct 10.0010.0010.00
4 Llama 3.1 8B Instruct 9.009.009.00
5 Command R 08-2024 7.007.007.00
6 OpenAI o1-mini 6.006.006.00
6 Hermes 3 Llama 3.1 70B 6.006.006.00
6 Mistral Nemo 6.006.006.00
9 GPT-4o-mini 4.004.004.00
9 DeepSeek V2.5 4.004.004.00
11 Claude 3.5 Sonnet 3.003.003.00
11 Qwen 2.5 72B Instruct 3.003.003.00
11 WizardLM-2 8x22B 3.003.003.00
14 Phi-3.5-mini Instruct 1.001.001.00
@@
1 Llama 3.1 70B Instruct 10.0810.0810.08
2 Gemini 1.5 Pro 002 9.249.249.24
3 Qwen 2.5 72B Instruct 5.885.885.88
4 Llama 3.1 405B Instruct 3.363.363.36
4 Llama 3.1 8B Instruct 3.363.363.36
6 Hermes 3 Llama 3.1 70B 2.522.522.52
7 OpenAI o1-mini 1.681.681.68
7 DeepSeek V2.5 1.681.681.68
7 Mistral Nemo 1.681.681.68
10 Command R 08-2024 0.840.840.84
10 Phi-3.5-mini Instruct 0.840.840.84
10 WizardLM-2 8x22B 0.840.840.84
13 Claude 3.5 Sonnet 0.000.000.00
13 GPT-4o-mini 0.000.000.00
@@
1 Llama 3.1 70B Instruct 12.4412.4412.44
2 Qwen 2.5 72B Instruct 8.818.818.81
3 OpenAI o1-mini 8.298.298.29
4 Gemini 1.5 Pro 002 7.777.777.77
5 Claude 3.5 Sonnet 6.746.746.74
5 Mistral Nemo 6.746.746.74
7 Llama 3.1 405B Instruct 6.226.226.22
7 GPT-4o-mini 6.226.226.22
9 DeepSeek V2.5 4.154.154.15
9 Llama 3.1 8B Instruct 4.154.154.15
9 Command R 08-2024 4.154.154.15
9 Phi-3.5-mini Instruct 4.154.154.15
9 WizardLM-2 8x22B 4.154.154.15
14 Hermes 3 Llama 3.1 70B 3.633.633.63
@@
1 Llama 3.1 70B Instruct 14.0014.0014.00
2 Llama 3.1 8B Instruct 10.0010.0010.00
3 Llama 3.1 405B Instruct 8.008.008.00
4 Hermes 3 Llama 3.1 70B 3.003.003.00
5 OpenAI o1-mini 2.022.022.02
6 Gemini 1.5 Pro 002 2.002.002.00
6 Mistral Nemo 2.002.002.00
8 Qwen 2.5 72B Instruct 1.001.001.00
8 Command R 08-2024 1.001.001.00
8 WizardLM-2 8x22B 1.001.001.00
11 Claude 3.5 Sonnet 0.000.000.00
11 DeepSeek V2.5 0.000.000.00
13 Phi-3.5-mini Instruct 1.00-1.00- 1.00
14 GPT-4o-mini 1.01-1.01- 1.01
@@
1 Llama 3.1 70B Instruct 24.0024.0024.00
2 Gemini 1.5 Pro 002 13.0013.0013.00
3 Llama 3.1 405B Instruct 11.0011.0011.00
4 Qwen 2.5 72B Instruct 6.006.006.00
4 Mistral Nemo 6.006.006.00
6 OpenAI o1-mini 5.005.005.00
7 Llama 3.1 8B Instruct 4.004.004.00
7 Hermes 3 Llama 3.1 70B 4.004.004.00
7 WizardLM-2 8x22B 4.004.004.00
10 DeepSeek V2.5 3.003.003.00
10 Command R 08-2024 3.003.003.00
12 Claude 3.5 Sonnet 2.002.002.00
12 Phi-3.5-mini Instruct 2.002.002.00
14 GPT-4o-mini 1.001.001.00
@@
1 Gemini 1.5 Pro 002 9.009.009.00
1 Llama 3.1 70B Instruct 9.009.009.00
3 Llama 3.1 8B Instruct 8.008.008.00
3 Mistral Nemo 8.008.008.00
5 Command R 08-2024 6.006.006.00
5 WizardLM-2 8x22B 6.006.006.00
7 OpenAI o1-mini 5.005.005.00
7 Claude 3.5 Sonnet 5.005.005.00
7 Hermes 3 Llama 3.1 70B 5.005.005.00
7 DeepSeek V2.5 5.005.005.00
11 Llama 3.1 405B Instruct 4.004.004.00
11 Qwen 2.5 72B Instruct 4.004.004.00
11 GPT-4o-mini 4.004.004.00
14 Phi-3.5-mini Instruct 1.001.001.00
@@
1 OpenAI o1-mini 9.009.009.00
2 Llama 3.1 405B Instruct 8.008.008.00
3 Claude 3.5 Sonnet 7.007.007.00
3 Gemini 1.5 Pro 002 7.007.007.00
3 WizardLM-2 8x22B 7.007.007.00
6 Llama 3.1 8B Instruct 4.004.004.00
6 Hermes 3 Llama 3.1 70B 4.004.004.00
8 Qwen 2.5 72B Instruct 3.003.003.00
8 Mistral Nemo 3.003.003.00
10 GPT-4o-mini 2.002.002.00
10 Llama 3.1 70B Instruct 2.002.002.00
10 DeepSeek V2.5 2.002.002.00
10 Command R 08-2024 2.002.002.00
14 Phi-3.5-mini Instruct 4.00-4.00- 4.00
@@
1 OpenAI o1-mini 8.608.608.60
2 Qwen 2.5 72B Instruct 6.996.996.99
3 Llama 3.1 70B Instruct 6.456.456.45
4 Llama 3.1 405B Instruct 5.385.385.38
5 Hermes 3 Llama 3.1 70B 4.844.844.84
6 Gemini 1.5 Pro 002 4.304.304.30
6 Llama 3.1 8B Instruct 4.304.304.30
8 Claude 3.5 Sonnet 3.763.763.76
9 Mistral Nemo 1.611.611.61
9 DeepSeek V2.5 1.611.611.61
9 Command R 08-2024 1.611.611.61
12 GPT-4o-mini 1.081.081.08
13 Phi-3.5-mini Instruct 0.540.540.54
14 WizardLM-2 8x22B 0.000.000.00
@@
1 Llama 3.1 70B Instruct 8.008.008.00
1 Llama 3.1 405B Instruct 8.008.008.00
3 WizardLM-2 8x22B 7.007.007.00
4 OpenAI o1-mini 6.006.006.00
4 Qwen 2.5 72B Instruct 6.006.006.00
6 Gemini 1.5 Pro 002 4.004.004.00
6 GPT-4o-mini 4.004.004.00
6 DeepSeek V2.5 4.004.004.00
6 Phi-3.5-mini Instruct 4.004.004.00
10 Claude 3.5 Sonnet 3.003.003.00
11 Command R 08-2024 2.002.002.00
11 Mistral Nemo 2.002.002.00
13 Hermes 3 Llama 3.1 70B 1.001.001.00
14 Llama 3.1 8B Instruct 1.00-1.00- 1.00
@@
1 Llama 3.1 70B Instruct 19.0019.0019.00
2 Llama 3.1 8B Instruct 12.0012.0012.00
3 Gemini 1.5 Pro 002 10.0010.0010.00
3 Llama 3.1 405B Instruct 10.0010.0010.00
5 Hermes 3 Llama 3.1 70B 9.009.009.00
6 OpenAI o1-mini 8.008.008.00
7 Mistral Nemo 7.007.007.00
8 Claude 3.5 Sonnet 6.006.006.00
8 WizardLM-2 8x22B 6.006.006.00
10 DeepSeek V2.5 5.005.005.00
11 Qwen 2.5 72B Instruct 4.004.004.00
11 Command R 08-2024 4.004.004.00
13 GPT-4o-mini 2.002.002.00
14 Phi-3.5-mini Instruct 1.001.001.00
@@
1 Hermes 3 Llama 3.1 70B 10.0010.0010.00
2 OpenAI o1-mini 9.009.009.00
2 DeepSeek V2.5 9.009.009.00
4 Gemini 1.5 Pro 002 8.008.008.00
4 Llama 3.1 405B Instruct 8.008.008.00
4 Qwen 2.5 72B Instruct 8.008.008.00
7 Llama 3.1 70B Instruct 6.006.006.00
8 WizardLM-2 8x22B 5.005.005.00
8 Command R 08-2024 5.005.005.00
10 Llama 3.1 8B Instruct 4.004.004.00
10 Mistral Nemo 4.004.004.00
10 Claude 3.5 Sonnet 4.004.004.00
10 Phi-3.5-mini Instruct 4.004.004.00
14 GPT-4o-mini 3.003.003.00